OpenStudy (anonymous):

REmainder and factor theorm 3x^2 + 6x - 4 = p(x-q)^2 + r

OpenStudy (anonymous):

compare this correspondingly..p(x^2)-2pq(x)+p*q^2+r with equation p=3 2pq=6 => hence 2*3*q=6 => q=1 as we have p=3,q=1 -4=p*q^2+r -4=3*1+r r=-7